OpenStudy (anonymous):

5/13 t = –9 Can someone tell me how to solve this step by step ?

OpenStudy (anonymous):

Isolate "t" by performing the same arithmetic on each side of the equal sign. So, if 5x - 3 = 17 : Add 3 to both sides to get: 5x - 3 [+ 3] = 17 [+ 3] simplified as 5x = 20 Then divide both sides by 5 to get: 5x/5 = 20/5 simplified as x = 4

OpenStudy (anonymous):

5/13t = -9 or 5t/13=-9, which one ?

OpenStudy (anonymous):

5/13t=-9 5=-117t ( multiplying both sides by 13t ) t=-( 5/117 )

OpenStudy (swag):

-( 5/117 )